owl shift 12/21/99

Theodoros Alexopoulos (theoalex@fnal.gov)
Tue, 21 Dec 1999 07:17:21 -0600 (CST)

Shift Summary for owl shift, 12/21/99

On Shift: ES, TA

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):7h (cosmics)

Accounting of lost time:

========================================================================
Runs ended this shift: None

Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
No beam. Tevatron recovery in progress. Best estimate for HEP is
late evening shift Tuesday according to the TV message.

========================================================================
Description of Shift:

run cosmics configuration.

========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift

o Main Alarm Screen terminal is not functioning; needs to be
replaced.